OpenStudy (anonymous):

Find the value of y when x equals 3 in this equation. -4x + 5y = -2 3 2 -2 -4 help anyone?

OpenStudy (anonymous):

OK, there are two ways to solve this so lets do the easy one

OpenStudy (anonymous):

Substitute x=3 into your x and then solve it for y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the answer 3?

OpenStudy (anonymous):

no, but close!

OpenStudy (anonymous):

-12+5y=-2 5y=-2+12

OpenStudy (anonymous):

hold up im doing something wrong..

OpenStudy (dannyo19):

If -12+5y=-2 we can add 12 to both sides

OpenStudy (dannyo19):

divide 12-2 by 5 and you get your answer.

OpenStudy (mathmale):

@teamworkk: Please show all of your work. The rest of us could give you more meaningful feedback if you'd do that.